# Quantum metrics from length functions on quantum groups

### v1

@@ 2.2. Quantum groups
In this section, we survey the relevant background results from the theory of compact quantum groups, following the $C^{*}$-algebraic approach initiated by Woronowicz [Wor98]; for detailed expositions, see [KS97] or [Tim08]). A compact quantum group is a pair $(A,\Delta)$ consisting of a unital $C^{*}$-algebra $A$ and a unital *-homomorphism $\Delta\colon A\to A\otimes A$ satisfying the coassociativity condition $(\Delta\otimes{\operatorname{id}})\Delta=({\operatorname{id}}\otimes\Delta)\Delta$ and such that the sets $\Delta(A)(1\otimes A)$ and $\Delta(A)(A\otimes 1)$ span dense subspaces in $A\otimes A$. Recall that here, and below,$\otimes$” is used to denote the minimal tensor product of $C^{*}$-algebras.
@@
Given a compact Hausdorff group $G$, setting $A=C(G)$ and letting $\Delta$ be the map dual to the group multiplication, one obtains a compact quantum group. Even better, all commutative examples take this form, making the theory of compact quantum groups perfectly compatible with Gelfand duality. Non-commutative examples may be constructed from a discrete group $\Gamma$, by setting $A=C^{*}_{\operatorname{red}}(\Gamma)$ and $\Delta(\lambda_{\gamma})=\lambda_{\gamma}\otimes\lambda_{\gamma}$, where $\lambda$ denotes the left regular representation of $\Gamma$. The quantum group literature contains numerous interesting examples not arising from compact or discrete groups, for instance Woronowicz’ quantum $SU(2)$ [Wor87] and Wang’s liberations of the unitary groups, the orthogonal groups and the permutation groups [Wan95, Wan98].
[…]
@@
    The Hilbert space is denoted by $L^{2}({\mathbb{G}})$ and the canonical map from $A$ to $L^{2}({\mathbb{G}})$ is denoted by $\Lambda$.
@@
    The GNS representation is denoted by $\lambda\colon A\to{\mathbb{B}}(L^{2}({\mathbb{G}}))$ and its image $\lambda(A)$ by $C({\mathbb{G}})$.
@@
    The von Neumann algebra generated by $C({\mathbb{G}})\subseteq{\mathbb{B}}(L^{2}({\mathbb{G}}))$ is denoted $L^{\infty}({\mathbb{G}})$.
The comultiplication descends to the $C^{*}$-algebra $C({\mathbb{G}})$, turning it into a compact quantum group in its own right, whose Haar state is the vector state $\lambda(a)\mapsto\left\langle\Lambda(1),\lambda(a)\Lambda(1)\right\rangle$. In what follows, we will only be concerned with this represented version of the quantum group, and therefore also denote the aforementioned vector state by $h$, the induced comultiplication by $\Delta$ and the canonical injection $C({\mathbb{G}})\to L^{2}({\mathbb{G}})$ by $\Lambda$. Whenever notationally convenient, we shall suppress the map $\Lambda$ and simply view $C({\mathbb{G}})$ as a subspace $L^{2}({\mathbb{G}})$.
@@ 2.2.1. The fundamental unitaries
The comultiplication therefore extends to a normal *-homomorphism $\Delta\colon L^{\infty}({\mathbb{G}})\to L^{\infty}({\mathbb{G}})\bar{\otimes}L^{\infty}({\mathbb{G}})$ (still implemented by $W$ and $V$) turning $L^{\infty}({\mathbb{G}})$ into a compact von Neumann algebraic quantum group [KV03, Tim08]. Following the established conventions in the literature, we think of $(C({\mathbb{G}}),\Delta)$ and $(L^{\infty}({\mathbb{G}}),\Delta)$ as different operator algebraic realizations of the (non-existing) underlying quantum group ${\mathbb{G}}$, which is indicated linguistically by phrases such as “Let ${\mathbb{G}}$ be a compact quantum group….”.
@@ 2.2.2. Flips and leg-numbering notation
for $a,b\in{\mathbb{B}}(L^{2}({\mathbb{G}}))$. We will also be using the standard leg-numbering notation, meaning that for an operator $T\in{\mathbb{B}}(L^{2}({\mathbb{G}})\hat{\otimes}L^{2}({\mathbb{G}}))$, $T_{12}$ denotes $T\otimes 1$, $T_{13}$ denotes $(1\otimes\sigma)(T\otimes 1)$, etc. Lastly, whenever $a\in C({\mathbb{G}})$ satisfies that $\Delta(a)\in C({\mathbb{G}})\odot C({\mathbb{G}})$ we will oftentimes use the Sweedler notation and write $\Delta(a)=a_{(1)}\otimes a_{(2)}$ for ease of notation. A similar remark applies to more general coactions of ${\mathbb{G}}$, to be introduced in Section 2.2.4 below.
@@ 2.2.3. Corepresentation theory
The corepresentation theory of ${\mathbb{G}}$ mirrors, in many ways, the representation theory of a classical compact group, in that one has natural notions of direct sums, tensor products, equivalence and irreducibility. Moreover, every finite-dimensional unitary corepresentation decomposes as a direct sum of irreducible ones. We let $I=\operatorname{Irred}({\mathbb{G}})$ denote the set of equivalence classes of irreducible corepresentations and let $(u^{\alpha})_{\alpha\in I}$ denote a fixed set of representatives for these equivalence classes. The unit in $C({\mathbb{G}})$ is a 1-dimensional (irreducible) corepresentation, and we denote its class in $\operatorname{Irred}({\mathbb{G}})$ by $e$. Each $u^{\alpha}$ is an element in $L^{\infty}({\mathbb{G}})\bar{\otimes}{\mathbb{B}}(H_{\alpha})$ for a finite-dimensional Hilbert space $H_{\alpha}$ (say, of dimension $d_{\alpha})$, and upon choosing an orthonormal basis for $H_{\alpha}$ we obtain an identification $L^{\infty}({\mathbb{G}})\bar{\otimes}{\mathbb{B}}(H_{\alpha})=\mathbb{M}_{d_{\alpha}}(L^{\infty}({\mathbb{G}}))$, and may thus also consider the associated matrix coefficients $u_{ij}^{\alpha}\in L^{\infty}({\mathbb{G}})$. It turns out, that $u_{ij}^{\alpha}\in C({\mathbb{G}})$ and that the set
@@
[…]
Tensor products and direct sums of corepresentations are denoted $\oplus$ and $\boxtimes$, respectively. For $\alpha\in\operatorname{Irred}({\mathbb{G}})$, the conjugate corepresentation is defined as $\overline{u^{\alpha}}:=((u_{ij}^{\alpha})^{*})$. This still satisfies the corepresentation relation (2.4) but may fail to be a unitary matrix — it is, however, equivalent to a unique element in $\operatorname{Irred}({\mathbb{G}})$ which we will denote by $\bar{\alpha}$; see e.g. [Tim08, Theorem 5.3.3 and Corollary 5.3.10]. Lastly, we denote by $N_{\alpha,\beta}^{\gamma}$ the multiplicity of $\gamma\in\operatorname{Irred}({\mathbb{G}})$ in the decomposition of $\alpha\boxtimes\beta$ into irreducibles; i.e.
@@
In the case of a classical compact group, the above constructions of course agree with their classical counterparts. For examples of the form $C^{*}_{\operatorname{red}}(\Gamma)$ all irreducible corepresentations are 1-dimensional and $(\lambda_{\gamma})_{\gamma\in\Gamma}$ is a complete set of representatives for $\operatorname{Irred}({\mathbb{G}})$, and $\operatorname{Pol}({\mathbb{G}})$ agrees with the group algebra ${\mathbb{C}}\Gamma\subset C^{*}_{\operatorname{red}}(\Gamma)$.
@@ 2.2.4. Coactions
A (left) coaction of ${\mathbb{G}}$ on a Hilbert space $H$ is a normal, unital, injective *-homomorphism $\delta\colon{\mathbb{B}}(H)\to L^{\infty}({\mathbb{G}})\overline{\otimes}{\mathbb{B}}(H)$ satisfying $({\operatorname{id}}\otimes\delta)\delta=(\Delta\otimes{\operatorname{id}})\delta$, see [Vae01]. For every unitary corepresentation $U\in L^{\infty}({\mathbb{G}})\overline{\otimes}{\mathbb{B}}(H)$, one obtains an associated coaction $\delta\colon{\mathbb{B}}(H)\to L^{\infty}({\mathbb{G}})\overline{\otimes}{\mathbb{B}}(H)$ by setting $\delta(T)=U^{*}(1\otimes T)U$.
@@ 2.2.5. Coamenability
A compact quantum group ${\mathbb{G}}$ is said to be coamenable, if the counit $\epsilon\colon\operatorname{Pol}({\mathbb{G}})\to\mathbb{C}$ extends to a character on $C({\mathbb{G}})$. In the case of a classical discrete group $\Gamma$, the counit $\epsilon\colon\mathbb{C}\Gamma\to\mathbb{C}$ is simply the extension of the trivial representation, and in this situation the notion of coamenability agrees with amenability of the group $\Gamma$, see e.g. [BO08]. The notion of coamenability has been studied extensively, and may be characterised in a number of different ways; see e.g. [Tom06, BMT01] for details.
Other examples of coamenable quantum groups include all commutative examples (where the counit is given by evaluation at the neutral element in the underlying compact group), Woronowicz’s quantum $SU(2)$ [Ban99a], and the quantum permutation group $S_{4}^{+}$ on four elements [Ban99b].
@@ 2.2.6. The fusion algebra
Letting $F({\mathbb{G}})$ denote the formal ${\mathbb{C}}$-vector space with basis $\operatorname{Irred}({\mathbb{G}})$, we obtain a *-algebra called the fusion algebra, in which the sum is induced by the direct summand, the involution is induced by conjugation $\alpha\mapsto\bar{\alpha}$ and the product is induced by (decomposition of) the tensor product of corepresentations:
@@ 2.2.7. The algebra of central functions
The character map $\chi\colon F({\mathbb{G}})\to\operatorname{Pol}_{z}({\mathbb{G}})$, given by $\chi(\alpha)=\sum_{i=1}^{d_{\alpha}}u^{\alpha}_{ii}$, extends by linearity to a unital *-isomorphism [Tim08, Proposition 3.2.14], and in this way we may view $F({\mathbb{G}})$ as a subalgebra of $\operatorname{Pol}({\mathbb{G}})$ whenever convenient. Moreover, it holds that $\operatorname{Pol}_{z}({\mathbb{G}})$ is dense in $C_{z}({\mathbb{G}})$ and $L_{z}^{\infty}({\mathbb{G}})$ in the norm- and strong operator topology, respectively, as was shown in [AC17]. One has an $h$-preserving conditional expectation $E\colon C({\mathbb{G}})\to C_{z}({\mathbb{G}})$ exactly when ${\mathbb{G}}$ is of Kac type (i.e. when $h$ is a trace) [Wan17, Lemma 6.3], and this conditional expectation will play a key role in the proof of Theorem A.
@@ 2.2.8. The dual
Associated with ${\mathbb{G}}$ is its dual discrete quantum group, $\hat{{\mathbb{G}}}$, [Tim08, Section 3.3], which also allows for a Hopf-algebraic, a $C^{*}$-algebraic and a von Neumann algebraic description. We denote the three algebras as
@@
The multiplicative unitaries, $W$ and $V$, introduced above are closely related to the dual pair $({\mathbb{G}},\hat{{\mathbb{G}}})$, in that
@@
The main focus in the present paper is on quantum metric structures on the $C^{*}$-algebra $C({\mathbb{G}})$. If $C({\mathbb{G}})$ admits the structure of a compact quantum metric space, then $C({\mathbb{G}})$ is necessarily separable, as is seen from the following standard argument: if the state space ${\mathcal{S}}(C({\mathbb{G}}))$ is metrised, it is itself separable and hence so is $C({\mathcal{S}}(C({\mathbb{G}})))$. The positive elements $C({\mathbb{G}})_{+}$ embed isometrically into $C({\mathcal{S}}(C({\mathbb{G}})))$ as evaluation functionals, so also $C({\mathbb{G}})_{+}$ is separable. Since every element in $C({\mathbb{G}})$ is a linear combination of four positive elements, $C({\mathbb{G}})$ is separable. We will therefore enforce separability of $C({\mathbb{G}})$ as a standing assumption throughout the rest of the paper, without further mention. Note that this implies that $L^{2}({\mathbb{G}})$ is a separable Hilbert space and that $\operatorname{Irred}({\mathbb{G}})$ is at most countable.

### v3

@@ 2.2. Quantum groups
[…]
@@
Given a compact Hausdorff group $G$, setting $A=C(G)$ and letting $\Delta$ be the map dual to the group multiplication, one obtains a compact quantum group. Even better, all commutative examples take this form, making the theory of compact quantum groups perfectly compatible with Gelfand duality. Non-commutative examples may be constructed from a discrete group $\Gamma$, by setting $A=C^{*}_{\operatorname{red}}(\Gamma)$ and $\Delta(\lambda_{\gamma})=\lambda_{\gamma}\otimes\lambda_{\gamma}$, where $\lambda$ denotes the left regular representation of $\Gamma$. The quantum group literature contains numerous interesting examples not arising from compact or discrete groups, for instance Woronowicz’ quantum $SU(2)$ [Wor:UAC] and Wang’s liberations of the unitary groups, the orthogonal groups and the permutation groups [Wang:free-products, Wang:quantum-symmetry].
One of the main features of the theory is the existence of a unique bi-invariant state $h\colon A\rightarrow$ $\mathbb{C}$}, meaning that
@@
    The Hilbert space is denoted by $L^{2}({\mathbb{G}})$, its norm by $\|\cdot\|_{2}$, and the canonical map from $A$ to $L^{2}({\mathbb{G}})$ is denoted by $\Lambda$.
@@
    The GNS representation is denoted by $\lambda\colon A\rightarrow{\mathbb{B}}(L^{2}({\mathbb{G}}))$ and its image $\lambda(A)$ by $C({\mathbb{G}})$.
@@
    The von Neumann algebra generated by $C({\mathbb{G}})\subseteq{\mathbb{B}}(L^{2}({\mathbb{G}}))$ is denoted by $L^{\infty}({\mathbb{G}})$.
The comultiplication descends to the $C^{*}$-algebra $C({\mathbb{G}})$, turning it into a compact quantum group in its own right, whose Haar state is the vector state $\lambda(a)\mapstochar\rightarrow\left\langle\Lambda(1),\lambda(a)\Lambda(1)\right\rangle$. In what follows, we will only be concerned with this represented version of the quantum group, and therefore also denote the aforementioned vector state by $h$, the induced comultiplication by $\Delta$ and the canonical injection $C({\mathbb{G}})\rightarrow L^{2}({\mathbb{G}})$ by $\Lambda$. Whenever notationally convenient, we shall suppress the map $\Lambda$ and simply view $C({\mathbb{G}})$ as a subspace of $L^{2}({\mathbb{G}})$.
@@ 2.2.1. The fundamental unitaries
The comultiplication therefore extends to a normal *-homomorphism $\Delta\colon L^{\infty}({\mathbb{G}})\rightarrow L^{\infty}({\mathbb{G}})\bar{\otimes}L^{\infty}({\mathbb{G}})$ (still implemented by $W$ and $V$) turning $L^{\infty}({\mathbb{G}})$ into a compact von Neumann algebraic quantum group [Kustermans-Vaes:vNA, Timmermann-book]. Following the established conventions in the literature, we think of $(C({\mathbb{G}}),\Delta)$ and $(L^{\infty}({\mathbb{G}}),\Delta)$ as different operator algebraic realizations of the (non-existing) underlying quantum group $\mathbb{G}$}, which is indicated linguistically by phrases such as “Let $\mathbb{G}$} be a compact quantum group….”.
@@ 2.2.2. Flips and leg-numbering notation
for $a,b\in{\mathbb{B}}(L^{2}({\mathbb{G}}))$. We will also be using the standard leg-numbering notation, meaning that for an operator $T\in{\mathbb{B}}(L^{2}({\mathbb{G}})\hat{\otimes}L^{2}({\mathbb{G}}))$, $T_{12}$ denotes $T\otimes 1$, $T_{13}$ denotes $(1\otimes\sigma)(T\otimes 1)$, etc. Lastly, whenever $a\in C({\mathbb{G}})$ satisfies that $\Delta(a)\in C({\mathbb{G}})\odot C({\mathbb{G}})$ we will frequently use the Sweedler notation and write $\Delta(a)=a_{(1)}\otimes a_{(2)}$ for ease of notation. A similar remark applies to more general coactions of $\mathbb{G}$}, to be introduced in Section 2.2.4 below.
@@ 2.2.3. Corepresentation theory
The corepresentation theory of $\mathbb{G}$} mirrors, in many ways, the representation theory of a classical compact group, in that one has natural notions of direct sums, tensor products, equivalence and irreducibility. Moreover, every finite-dimensional unitary corepresentation decomposes as a direct sum of irreducible ones. We let $I=\operatorname{Irred}({\mathbb{G}})$ denote the set of equivalence classes of irreducible corepresentations and let $(u^{\alpha})_{\alpha\in I}$ denote a fixed set of representatives for these equivalence classes. The unit in $C({\mathbb{G}})$ is a 1-dimensional (irreducible) corepresentation, and we denote its class in $\operatorname{Irred}({\mathbb{G}})$ by $e$. Each $u^{\alpha}$ is an element in $L^{\infty}({\mathbb{G}})\bar{\otimes}{\mathbb{B}}(H_{\alpha})$ for a finite-dimensional Hilbert space $H_{\alpha}$ (say, of dimension $d_{\alpha})$, and upon choosing an orthonormal basis for $H_{\alpha}$ we obtain an identification $L^{\infty}({\mathbb{G}})\bar{\otimes}{\mathbb{B}}(H_{\alpha})=\mathbb{M}_{d_{\alpha}}(L^{\infty}({\mathbb{G}}))$, and may thus also consider the associated matrix coefficients $u_{ij}^{\alpha}\in L^{\infty}({\mathbb{G}})$. It turns out that $u_{ij}^{\alpha}\in C({\mathbb{G}})$ and that the set
@@
forms a dense Hopf *-subalgebra in $C({\mathbb{G}})$. More precisely, $\Delta$ restricts to a comultiplication $\Delta\colon \operatorname{Pol}({\mathbb{G}})\rightarrow\operatorname{Pol}({\mathbb{G}})\odot\operatorname{Pol}({\mathbb{G}})$ given by $\Delta(u_{ij}^{\alpha})=\sum_{k=1}^{d_{\alpha}}u_{ik}^{\alpha}\otimes u_{kj}^{\alpha}$, and $\operatorname{Pol}({\mathbb{G}})$ can be further equipped with an antipode $S\colon\operatorname{Pol}({\mathbb{G}})\rightarrow\operatorname{Pol}({\mathbb{G}})$ and a counit $\epsilon\colon\operatorname{Pol}({\mathbb{G}})\rightarrow{\mathbb{C}}$; [Timmermann-book, Theorem 5.4.1]. For future reference, we note that since $\operatorname{Pol}({\mathbb{G}})$ is a Hopf *-algebra, $W^{*}$ and $V$ actually map $\operatorname{Pol}({\mathbb{G}})\odot\operatorname{Pol}({\mathbb{G}})$ bijectively onto itself (see e.g. [Timmermann-book, Theorem 1.3.18]) and the same is therefore true for $W$ and $V^{*}$.
Direct sums and tensor products of corepresentations are denoted $\oplus$ and $\boxtimes$, respectively. For $\alpha\in\operatorname{Irred}({\mathbb{G}})$, the conjugate corepresentation is defined as $\overline{u^{\alpha}}:=((u_{ij}^{\alpha})^{*})$. This still satisfies the corepresentation relation (2.4) but may fail to be a unitary matrix — it is, however, equivalent to a unique element in $\operatorname{Irred}({\mathbb{G}})$ which we will denote by $\bar{\alpha}$; see e.g. [Timmermann-book, Theorem 5.3.3 and Corollary 5.3.10]. Lastly, we denote by $N_{\alpha,\beta}^{\gamma}$ the multiplicity of $\gamma\in\operatorname{Irred}({\mathbb{G}})$ in the decomposition of $\alpha\boxtimes\beta$ into irreducibles; i.e.
@@
In the case of a classical compact group, the above constructions of course agree with their classical counterparts. For examples of the form $C^{*}_{\operatorname{red}}(\Gamma)$, all irreducible corepresentations are 1-dimensional and $(\lambda_{\gamma})_{\gamma\in\Gamma}$ is a complete set of representatives for $\operatorname{Irred}({\mathbb{G}})$, and $\operatorname{Pol}({\mathbb{G}})$ agrees with the group algebra ${\mathbb{C}}\Gamma\subset C^{*}_{\operatorname{red}}(\Gamma)$.
@@ 2.2.4. Coactions
A (left) coaction of $\mathbb{G}$} on a Hilbert space $H$ is a normal, unital, injective *-homomorphism $\delta\colon{\mathbb{B}}(H)\rightarrow L^{\infty}({\mathbb{G}})\bar{\otimes}{\mathbb{B}}(H)$ satisfying $({\operatorname{id}}\otimes\delta)\delta=(\Delta\otimes{\operatorname{id}})\delta$, see [Vaes:implementation]. For every unitary corepresentation $U\in L^{\infty}({\mathbb{G}})\bar{\otimes}{\mathbb{B}}(H)$, one obtains an associated coaction $\delta\colon{\mathbb{B}}(H)\rightarrow L^{\infty}({\mathbb{G}})\bar{\otimes}{\mathbb{B}}(H)$ by setting $\delta(T)=U^{*}(1\otimes T)U$.
@@ 2.2.5. Coamenability
A compact quantum group $\mathbb{G}$} is said to be coamenable, if the counit $\epsilon\colon\operatorname{Pol}({\mathbb{G}})\rightarrow\mathbb{C}$ extends to a character on $C({\mathbb{G}})$. In the case of a classical discrete group $\Gamma$, the counit $\epsilon\colon\mathbb{C}\Gamma\rightarrow\mathbb{C}$ is simply the extension of the trivial representation, so in this situation the notion of coamenability agrees with amenability of the group $\Gamma$, see e.g. [Brown-Ozawa]. The notion of coamenability has been studied extensively, and may be characterised in a number of different ways; see e.g. [tomatsu-amenable, bedos-murphy-tuset] for details.
Other examples of coamenable quantum groups include all commutative examples (where the counit is given by evaluation at the neutral element in the underlying compact group), Woronowicz’s quantum $SU(2)$ [banica-subfactor], and the quantum permutation group $S_{4}^{+}$ on four elements [banica-sym].
@@ 2.2.6. The fusion algebra
Letting $F({\mathbb{G}})$ denote the formal $\mathbb{C}$}-vector space with basis $\operatorname{Irred}({\mathbb{G}})$, we obtain a *-algebra called the fusion algebra, in which the sum is induced by the direct sum, the involution is induced by conjugation $\alpha\mapstochar\rightarrow\bar{\alpha}$ and the product is induced by (decomposition of) the tensor product of corepresentations:
@@
Remark 2.10.
Note that if two quantum groups ${\mathbb{G}}_{1}$ and ${\mathbb{G}}_{2}$ are monoidally equivalent (see e.g. [NeshveyevTusetBook, Definition 2.3.8]), then the equivalence induces a bijection between $\operatorname{Irred}({\mathbb{G}}_{1})$ and $\operatorname{Irred}({\mathbb{G}}_{2})$ in such a way that the vector space isomorphism $F({\mathbb{G}}_{1})\cong F({\mathbb{G}}_{2})$ is a *-homomorphism.
@@ 2.2.7. The algebra of central functions
The character map $\chi\colon F({\mathbb{G}})\rightarrow\operatorname{Pol}_{z}({\mathbb{G}})$, given by $\chi(\alpha)=\sum_{i=1}^{d_{\alpha}}u^{\alpha}_{ii}$, extends by linearity to a unital *-isomorphism [BanicaIntro2022, Proposition 3.23 and Theorem 3.24], and in this way we may view $F({\mathbb{G}})$ as a subalgebra of $\operatorname{Pol}({\mathbb{G}})$ whenever convenient. Moreover, it holds that $\operatorname{Pol}_{z}({\mathbb{G}})$ is dense in $C_{z}({\mathbb{G}})$ and $L_{z}^{\infty}({\mathbb{G}})$ in the norm- and strong operator topology, respectively, as was shown in [Crann:CaracterDensity]. One has an $h$-preserving conditional expectation $E\colon C({\mathbb{G}})\rightarrow C_{z}({\mathbb{G}})$ exactly when $\mathbb{G}$} is of Kac type (i.e. when $h$ is a trace) [Wang:lacunary, Lemma 6.3], and this conditional expectation will play a key role in the proof of Theorem A.
@@ 2.2.8. The dual
Associated with $\mathbb{G}$} is its dual discrete quantum group, $\hat{{\mathbb{G}}}$, [Timmermann-book, Section 3.3], which also allows for a Hopf-algebraic, a $C^{*}$-algebraic and a von Neumann algebraic description. We denote the three algebras as
@@
[…]
@@
The main focus in the present paper is on quantum metric structures on the $C^{*}$-algebra $C({\mathbb{G}})$. If $C({\mathbb{G}})$ admits the structure of a compact quantum metric space, then $C({\mathbb{G}})$ is necessarily separable, as is seen from the following standard argument: if the state space ${\mathcal{S}}(C({\mathbb{G}}))$ is metrized, it is itself separable and hence so is $C({\mathcal{S}}(C({\mathbb{G}})))$. The positive elements $C({\mathbb{G}})_{+}$ embed isometrically into $C({\mathcal{S}}(C({\mathbb{G}})))$ as evaluation functionals, so also $C({\mathbb{G}})_{+}$ is separable. Since every element in $C({\mathbb{G}})$ is a linear combination of four positive elements, $C({\mathbb{G}})$ is separable. We will therefore enforce separability of $C({\mathbb{G}})$ as a standing assumption throughout the rest of the paper, without further mention. Note that this implies that $L^{2}({\mathbb{G}})$ is a separable Hilbert space and that $\operatorname{Irred}({\mathbb{G}})$ is at most countable.